OpenStudy (anonymous):

Anyone know what the method of elimination is??? Thx also, this a bit random but What imression do we get of scrooge when it says he is "hard and sharp as flint" in a christmas carol?? :)

OpenStudy (anonymous):

elimination method or the addition or subtraction method. In this way of solving systems of equations, one variable is eliminated by adding or subtracting the equations (understand their names now? =P).

OpenStudy (anonymous):

Put one equation on top of the other and solve for one variable. When adding two equations, you basically add all parts of them. Say you had the equations: 4x + 5y = 14 -4x - 3y = -10 Adding them would give 2y = 4

OpenStudy (anonymous):

Ohhhh hahaha thanks a lot!! :p